# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ II.2 Planar geometries
Single-plane crystals of ions have been prepared in both Penning and rf ion traps by adjusting the confinement to be much stronger in one direction relative to the two orthogonal directions, see Fig. 21 for some representative examples. In contrast with the work on linear 1D ion crystals, developments such as near ground state cooling of the transverse motion in these crystals has only been demonstrated in the last $\sim 5$ years [162, 246, 171, 129]. We note that 2D crystals of ions are at the basis of protocols for realising robust quantum information processing [311, 11]. They have been employed in quantum simulation experiments of long-range Ising spin models, producing entanglement of the internal spin degrees of freedom [27, 118], and in the preparation of geometrically frustrated ground states [129, 259]. Two-dimensional ion crystals have also been used to simulate spin-boson models [273, 124].
In this section, after reviewing the different trap configurations for preparing 2D ion crystals, we discuss their thermodynamic (ground state configuration) and dynamic (normal mode) properties, and the structural instability when a single plane crystal transitions to multiple planes.
@@ II.2.1 Trap configurations
where $(x,y,z)$ denote coordinates along the principal axes of the confining potential. Harmonic confinement is assumed, which is typically a good approximation. For a Penning ion trap (see Sec. I.2.2), $(x,y,z)$ are coordinates in the rotating frame of the crystal with $z$ denoting distance along the magnetic field axis. For a linear rf trap (see Sec. I.2.1), $V$ denotes the confining pseudopotential and $z$ the distance along the nodal line in the center of the trap that is free from rf fields. As briefly mentioned in the above introductory paragraph, single plane ion crystals can be prepared by adjusting the confinement to be much stronger in one direction relative to the two orthogonal directions. Single plane crystals have been prepared in a variety of trap configurations, which we summarize below.
@@
In Penning traps, single-plane ion crystals are formed in a plane perpendicular to the magnetic field of the trap by adjusting $\beta_{x},\,\beta_{y}\ll 1$. This is achieved by controlling the ion crystal rotation frequency to produce a weak radial confinement compared to the axial confinement (see Fig. 7). The relative strength of $\beta_{x}$ and $\beta_{y}$ is set by the strength of the rotating wall electric field. Typically $\beta_{x}\sim\beta_{y}$, giving rise to ion crystal shapes that are approximately circular. Single-plane crystals of up to 400 to 500 ions have been produced and controlled in Penning traps [214, 10, 203]. Readout of the spin state of an individual ion in a single experiment [334, 205] and addressing of individual ions [205] in the rotating frame of the crystal has been demonstrated.
@@
Similar to the Penning trap, single-plane crystals have been prepared in rf traps by adjusting the confinement in the axial (z^)($\hat${z}) direction to be strong compared to the confinement in the radial direction [88, 263, 80, 166]. With traps operated in this configuration, which we denote the radial configuration, single-plane crystals of up to $\sim$50 ions have been prepared and well characterized. Another approach for forming 2D ion crystals is to weaken the confinement in a lateral direction to the axis ($\hat{z}-$direction) in a linear rf trap [25, 167, 20, 305, 330, 259, 171]. Starting with a 1D ion chain, weakening the lateral confinement in the $\hat{y}-$direction results in a 2D zig-zag structure as discussed in Sec. II.1.4. Through further weakening of the lateral confinement and careful control of the relative confinement strengths in the lateral and axial directions, single plane ion crystals that have an approximate elliptical shape are obtained with $1<\beta_{y}\ll\beta_{x}$. See Figs. 21 (b) and (c). With traps operated in this lateral configuration, single plane crystals of up to 150 ions have been observed in a surface electrode trap [305]. Very stable crystal structures have been demonstrated and studied with up to 105 ions in a three-layer rf trap [171] operated at room temperature and with up to 512 ions in a monolithic 3-layer trap operated at cryogenic temperatures [129].
@@
In contrast to 1D chains of ions in the linear rf trap, 2D single plane crystals in rf traps undergo substantial micromotion. In both the radial and lateral configurations discussed above, the driven micromotion of the ions at the rf drive frequency can be designed to lie in the plane of the ion crystal [263, 171]. A nice feature of the lateral configuration is that the trap can be designed so that the micromotion occurs in a single direction [171], which enables micromotion free interactions with laser beams from more than one direction.
@@ II.2.2 Ground state configuration
The ground state equilibrium configuration of a single-plane, trapped ion crystal is an approximate triangular lattice [91, 94, 278], like that shown in Fig. 22. In the interior of the ion crystal, the ions arrange themselves at the vertices of equilateral triangles, characteristic of an infinite 2D system. Such triangular lattices have frequently been called “hexagonal planes” in the literature as it emphasizes the 6-fold symmetry of the lattice.
Figure 22: Charge distribution in the ground state of a two-dimensional ion crystal. The dots correspond to the local minimum energy configuration evaluated numerically for $N=500$ identical charges in (a) a purely quadratic trap potential and (b) a quadratic trap potential plus a lowest-order anharmonic moment $(C_{4}\neq 0)$ designed to generate a more uniform lattice in the interior of the crystal. The red dots indicate dislocations in the crystal lattice. From 96.
For a harmonic trap potential such as Eq. (54), the 2D number density of a single plane crystal is inhomogeneous, decreasing from the center to the boundary of the crystal. For simplicity, consider the azimuthally symmetric confinement potential (Eq. (22) in Sec. I.2.2)
@@
obtained with $\beta_{x}=\beta_{y}\equiv\beta$. Single plane ion crystals are formed in the $\hat{z}=0$ plane of the trap when $\beta\ll 1$. For a single plane crystal of $N$ ions, $\beta$ must be less than a critical $\beta_{c}(N)$, $\beta<\beta_{c}(N)\approx 0.665/\sqrt{N}$ [94]. We can estimate the radial dependence in the 2D number density of a single plane crystal by considering three dimensional crystals formed in the potential of Eq. (55). This potential produces a crystal with constant, three-dimensional density $n_{0}$ (see Eqs. (19, 23)) given by
$n_{0}=\frac{\epsilon_{0}m}{Q^{2}}\omega_{z}^{2}(1+2\beta)$ (56)
in a spheroidal volume of axial extent $2Z_{p}$ and diameter $2R_{p}$. For this spheroidal shaped crystal, the 2D number density per unit area, $\sigma(\rho)$, is
@@
where $\sigma=n_{0}2Z_{p}$ is the 2D number density at the center. Equation (57) provides a good estimate for the 2D number density of a single plane crystal, obtained in the limit of small $Z_{p}/R_{p}$. It is the equivalent expression for the inhomogeneity in the 2D number density for a single plane crystal as Eq. (28) is for the inhomogeneity in the linear density for 1D chains of ions.
@@
For finite planes, the inhomogeneous charge density of single-plane ion crystals is necessarily associated with dislocations. These are illustrated by the red dots in Fig. 22(a) [96]. Near the boundary of the planar crystal, shells are formed that conform to the overall symmetry of the confining potential [90]. For azimutally symmetric potentials, the shells are circular and the locations of ions in neighboring shells is staggered, see Fig. 22. This shell structure has a different symmetry than the underlying triangular lattice and contributes to the formation of dislocations.
The number of dislocations in the interior of the crystal can be reduced by modifying the confining potential through the addition of anharmonic terms, thereby approaching a regular triangular lattice. The configuration of Fig. 22(a), for instance, can be made more regular by adding a fourth-order term to the harmonic confining potential. The resulting structure is presented in Fig. 22(b) for an optimized potential and in a configuration relevant for the Penning ion trap. The fourth-order term is frequently called a $C_{4}$ term. Structures even closer to uniform triangular lattices can be achieved by adding corrections with the same underlying symmetry as the target structure. In Penning ion traps the symmetry of the boundary is determined by the symmetry of the rotating wall potential. Theoretical studies indicate that a nearly perfect triangular lattice can be obtained by fine tuning the strength of an $m=3$ rotating wall potential [96], where $m$ refers to the azimuthal symmetry of the applied rotating wall potential ($\propto\cos[m(\phi+\omega_{r}t)]$, see Eq. (25)). Experimental work demonstrated that this strategy leads to the formation of more regular ion crystals in Penning traps [205]. Overall, these strategies are at early stages of experimental investigation.
We emphasize that the discussion in this section neglects micromotion and assumes the validity of the pseudopotential approximation in linear rf traps.
@@ II.2.3 Normal modes
In what follows we discuss the normal modes of single-plane ion crystals. We first neglect micromotion and assume that the ions are confined to a single plane by a harmonic, conservative potential.
@@
We start with the normal modes of an infinite, single plane ion crystal. This is an idealization, that provides intuition for understanding the normal modes of finite-sized, two-dimensional structures. In the infinite triangular lattice, the equilibrium positions are $\bm{\rho}=[(n_{1}+n_{2}/2)\hat{x}+(\sqrt{3}/2)\,n_{2}\,\hat{y}]a_{\mathrm{lat}}$, with $n_{1},n_{2}$ integers. The normal modes are waves with wave vector $\mathbf{k}$} in the xyx-y plane and displacements proportional to $\exp\left(i{\mathbf{k}}\cdot\bm{\rho}\right)$. For every $\mathbf{k}$} there are three different polarizations: two are in the xyx-y plane while the third is in the $\hat{z}-$direction, perpendicular to the plane. As an example consider the two in-plane modes when $k_{y}=0$. They have polarizations in the $\hat{x}$ and $\hat{y}$ directions, corresponding to compressional and transverse modes. Their dispersion relation, along with the dispersion relation for oscillations normal to the plane, is displayed in Fig. 23 for $\sigma a_{0}^{2}=1.07$ and at the instability of the single plane ($\sigma a_{0}^{2}=1.11$). Note that the bandwidth of the in-plane modes overlaps with the bandwidth of the mode for oscillations normal to the plane.
@@
The dispersion relation of the in-plane modes exhibits the characteristic features of acoustic modes in solids: the frequency increases as the wavelength decreases. The longitudinal (compressional) mode displays a sublinear dispersion at long wavelengths ($\hat{x}-$mode in Fig. 23), $\omega(k)\sim\sqrt{k}$, arising from the long-range nature of the Coulomb repulsion [303]. The mode corresponding to out-of-plane oscillations in the $\hat{z}$ direction is analogous to a drumhead mode. Its dispersion relation is similar to optical modes in solids: the mode frequency decreases as the wavelength increases. The largest frequency is the transverse confining frequency $\omega_{z}$, where the crystal oscillates as a bulk. The shortest wavelengths, comparable to the inter-ion spacing, are the smallest frequencies. As visible in Fig. 23, at the single-plane stability limit, $\left(\sigma{a_{0}}^{2}=1.1\right)$, the mode frequency becomes zero at $k_{x}a_{\mathrm{lat}}=4\pi/3$, producing a second order phase transition discussed in Sec. II.2.4.
@@
Figure 23: Normal mode frequencies of a single plane triangular lattice as a function of $k_{x}$, in units of the inverse of the lattice constant $a_{\mathrm{lat}}=(2/\sqrt{3}\sigma)^{1/2}$. The frequencies $\omega({\mathbf{k}})$ are squared and in units of the frequency $\omega_{z}$ of the transverse harmonic trap, confining the ions on the xyx-y plane. The modes shown are the in-plane (along $\hat{x}$ and $\hat{y}$) and the transverse (along $\hat{z}$) modes for $k_{y}=0$. Dotted lines: $\sigma a_{0}^{2}=1.07$; solid lines $\sigma a_{0}^{2}=1.11$, corresponding approximately to the density $\sigma$ where the single plane becomes mechanically unstable. The branch of the transverse modes approaches 0 for $k_{x}=4\pi/3$. The phase transition of this mechanical instability is discussed in Sec. II.2.4. From 94.
@@
The normal modes of trapped, single-plane ion crystals take on the character of the infinite single-plane crystal modes, including a similar dispersion relation. The general theoretical approach consists of diagonalizing the stiffness matrix, obtained by performing the Taylor expansion of the Coulomb potential, with some salient differences between the rf- and the Penning trap due to the respective micromotion and the strong magnetic field [329, 291]. In rf traps, some of the ions in the crystal undergo significant micromotion: Including the effects of micromotion in describing normal modes is challenging and has only been incorporated in Refs. [167, 171]. In Ref. [167] significant discrepancies with pseudopotential theory were observed with the measured in-plane mode spectrum of a three-ion crystal. Including micromotion through a full time-dependent Coulomb theory resulted in good agreement between the measured and predicted eigenfrequencies.
@@
In a Penning ion trap, the natural frame in which to analyze the normal modes is the frame that co-rotates with the ion crystal. Because the non-thermal ion motion is a rigid body rotation about the trap symmetry axis, it does not impact the calculation of the normal modes beyond adding a centrifugal potential. However, the strong magnetic field does modify the theoretical procedure for determining the ion crystal normal modes.
@@
In general, a single-plane crystal of $N$ ions supports $2N$ in-plane modes and $N$ transverse modes. The latter are frequently called drumhead modes because they resemble the modes of a drum with open boundary conditions [329, 277, 291]. In a Penning ion trap, the drumhead modes for a single plane crystal involve ion motion parallel to the magnetic field. Therefore, the magnetic field does not impact the drumhead mode frequencies or eigenvectors in the harmonic approximation, and the drumhead mode structure is the same for single-plane crystals in Penning traps and in rf traps, provided the pseudopotential approximation applies to the latter. In Ref. [171] reasonable agreement with the predictions of the pseudopotential model was obtained with the measured spectrum of the transverse modes of an 8-ion crystal generated in a rf trap. Theoretical calculations including the micromotion did not appear to improve the agreement with the experimental data.
@@
Figure 24 illustrates the oscillation amplitudes of the four highest-frequency drumhead modes for a single-plane ion crystal. From larger to lower frequency: The highest frequency mode is the center-of-mass (or bulk) mode where all ions in the crystal oscillate as a bulk at frequency $\omega_{z}$. The next drumhead modes are tilt modes. As the normal mode frequency further decreases, so does the effective wavelength. In the short wave-length limit, the wavelength is of the order of the lattice spacing in the interior of the crystal and only ions in the crystal interior participate in the mode. In this spatial region the crystal approaches a regular triangular lattice and the short-wavelength drumhead modes essentially coincide with the transverse modes of the infinite crystal.
@@
Figure 24: The four highest frequency drumhead mode eigenvectors calculated for a single plane ion crystal consisting of $N=217$ ions. The three rows correspond to three different in-plane confinement strengths, respectively. The color code illustrates the eigenvector amplitude with the maximal value set to one. From 329.
@@
Figure 25: Example calculation of the in-plane mode frequencies for a single plane crystal with $N=120$ ions in the presence of no magnetic field (green central branch) and in the presence of a strong magnetic field that splits the single green branch into low-frequency (blue) and high frequency (orange) branches. The green branch is the in-plane mode spectrum in the pseudopotential approximation for a single plane ion crystal in an rf trap with $\omega_{z}/(2\pi)=1.59$ MHz and effective radial confinement $\sqrt{\beta}\omega_{z}/(2\pi)=267.5$ kHz. The blue and orange branches are the low-frequency ${\mathbf{E}}\times{\mathbf{B}}$ mode and high-frequency cyclotron mode spectra for a single plane crystal in a Penning trap with identical axial ($\omega_{z}$) and effective radial ($\sqrt{\beta}\omega_{z}$) confinement. For the Penning trap, a cyclotron frequency $\Omega_{c}/(2\pi)=7.60$ MHz and a rotation frequency $\omega_{r}/(2\pi)=180$ kHz are assumed, and the mode frequencies are referenced to the rotating frame of the ion crystal. The calculation assumes a weak asymmetry in the radial confinement characterized by $\omega_{W}=68$ kHz (see Eq. (26)). From 291.
@@
For single-plane crystals in an rf trap where the magnetic field is nominally zero, the in-plane modes exhibit acoustic-like dispersion, similar to the in-plane modes (the $\hat{x}$ and $\hat{y}$ polarizations of Fig. 23) of an infinite crystal. The green branch in Fig. 25 shows the results of a calculation of the in-plane mode spectrum for a single-plane crystal consisting of 120 ions in zero magnetic field [291]. The calculation is made for an rf trap in the pseudopotential approximation with an axial frequency $\omega_{z}/(2\pi)=1.59$ MHz and trap anisotropy parameter $\beta=0.028$. For $N=120$ ions, there exist 240 in-plane modes, which are ordered in the figure according to increasing frequency. The lowest frequency modes are associated with the weak asymmetry that provides azimuthal confinement, without which there would be a zero-frequency mode corresponding to rigid body rotation about the $\hat{z}-$axis. The higher frequency modes correspond to effective shorter wavelength oscillations where ion motion becomes anti-correlated over a distance of a few inter-ion spacings.
@@
Adding a strong magnetic field normal to the single-plane crystal (parallel to the $z$-axis) gives rise to velocity dependent forces. This is the case of the Penning trap. Here the mode analysis becomes more involved for the in-plane modes, requiring the solution of a generalized eigenvalue problem [97, 291, 329]. Figure 25 illustrates that an important consequence of the strong magnetic field is to split the single band of $2N$ in-plane modes for $B=0$ (green central branch) into a band of $N$ high frequency cyclotron modes and a band of $N$ low frequency ${\mathbf{E}}\times{\mathbf{B}}$ modes with frequency on the order of the single ion magnetron frequency $\omega_{m}$ defined in Sec. I.2.2.

### v2

@@ II.2 Planar geometries
Single-plane crystals of ions have been prepared in both Penning and rf ion traps by adjusting the confinement to be much stronger in one direction relative to the two orthogonal directions. Figure 21 shows some representative examples. In contrast with ion chains, near ground state cooling of the transverse motion of single-plane crystals has only been demonstrated very recently [175, 186, 140, 273]. We note that two-dimensional ion crystals are at the basis of protocols for realizing robust quantum information processing [346, 15]. They have been employed in quantum simulation experiments of long-range Ising spin models, producing entanglement of the internal spin degrees of freedom [30, 128], and in the preparation of geometrically frustrated ground states [290, 140]. Two-dimensional ion crystals have also been used to simulate spin-boson models [305, 134].
In this section, after reviewing the different trap configurations for preparing 2D ion crystals, we discuss their thermodynamic (ground state configuration) and dynamic (normal mode) properties, and the structural instability when a single-plane crystal transitions to multiple planes. Many of the results of this section can be applied to single-plane crystals in rf traps, in the limit in which the pseudopotential approximation holds and micromotion can be neglected.
@@ II.2.1 Trap configurations
where $(x,y,z)$ denote coordinates along the principal axes of the confining potential and relative to the center of the trap. Harmonic confinement is assumed, which is typically a good approximation. For a Penning trap (see Sec. I.2.2), $(x,y,z)$ are coordinates in the rotating frame of the crystal with $z$ denoting distance along the magnetic field axis. For a linear rf trap, $V$ denotes the confining pseudopotential and $z$ the distance along the nodal line in the center of the trap that is free from rf fields (see Sec. I.2.1). Single-plane crystals have been prepared in a variety of trap configurations, which we summarize below.
@@
In Penning traps, single-plane ion crystals are formed in the plane perpendicular to the magnetic field of the trap by adjusting $\beta_{x},\,\beta_{y}\ll 1$. This is achieved by controlling the crystal rotation frequency to produce a weak radial confinement compared to the axial confinement (see Fig. 7). The relative strength of $\beta_{x}$ and $\beta_{y}$ is set by the strength of the rotating wall electric field. Typically $\beta_{x}\sim\beta_{y}$, giving rise to crystal shapes that are approximately circular. Single-plane crystals of up to 400 to 500 ions have been produced and controlled in Penning traps [238, 226, 11]. Readout of the spin state of an individual ion in a single experiment [229, 371] and addressing of individual ions [229] in the rotating frame of the crystal have been demonstrated.
@@
Similar to the Penning trap, single-plane crystals have been prepared in rf traps by adjusting the confinement in the axial (z^)($\hat${z}) direction to be strong compared to the confinement in the radial direction [93, 294, 86, 181]. With traps operated in this configuration (“radial configuration”), single-plane crystals of up to $\sim$50 ions have been prepared and well characterized. Another approach for forming 2D ion crystals is to weaken the confinement in a direction transverse to the axis of a linear rf trap [28, 23, 182, 340, 366, 290, 186]. Starting with an ion chain along the $\hat{z}-$direction, weakening the lateral confinement in the $\hat{y}-$direction results in a 2D zig-zag structure as discussed in Sec. II.1.4. By further weakening the lateral confinement and careful controlling of the relative confinement strengths in the transverse and axial directions to ensure $1<\beta_{y}\ll\beta_{x}$, single-plane ion crystals are obtained that have an approximate elliptical shape, see Figs. 21(b) and (c). With traps operated in this lateral configuration, single-plane crystals of up to 150 ions have been observed in a surface electrode trap [340]. Stable crystal structures have been realized and studied with up to 105 ions in a three-layer rf trap operated at room temperature [186] and with up to 512 ions in a monolithic 3-layer trap operated at cryogenic temperatures [140].
@@
In contrast to chains of ions in the linear rf trap, single-plane crystals in rf traps undergo substantial micromotion. In both the radial and lateral configurations discussed above, the driven micromotion of the ions at the rf drive frequency can be designed to lie in the plane of the ion crystal [294, 186]. A nice feature of the lateral configuration is that the trap can be designed so that the micromotion occurs in a single direction [186], which enables micromotion-free interactions with laser beams from more than one direction.
@@ II.2.2 Ground state configuration
The ground state equilibrium configuration of a single-plane, trapped ion crystal is an approximate triangular lattice [97, 100, 311], like the one shown in Fig. 22. In the interior of the ion crystal, the ions arrange themselves at the vertices of equilateral triangles, characteristic of an infinite 2D system. Such triangular lattices have frequently been called “hexagonal planes” in the literature as it emphasizes the 6-fold symmetry of the lattice.
Figure 22: (Color online) Charge distribution in the ground state of a two-dimensional ion crystal. The dots correspond to the local minimum energy configuration evaluated numerically for $N=500$ identical charges in (a) a purely quadratic trap potential and (b) a quadratic trap potential plus a lowest-order anharmonic moment $(C_{4}\neq 0)$ designed to generate a more uniform lattice in the interior of the crystal. The red (enlarged) dots indicate dislocations in the crystal lattice. From 102.
For a harmonic trap potential such as Eq. (54), the 2D number density of a single-plane crystal is inhomogeneous, decreasing from the center to the boundary of the crystal. For simplicity, consider the azimuthally symmetric confinement potential (Eq. (22) in Sec. I.2.2)
@@
obtained with $\beta_{x}=\beta_{y}\equiv\beta$. single-plane ion crystals are formed in the $\hat{z}=0$ plane of the trap when $\beta\ll 1$. For a single-plane crystal of $N$ ions, $\beta$ must be less than a critical $\beta_{c}(N)\approx 0.665/\sqrt{N}$ [100]. We can estimate the radial dependence in the 2D number density of a single-plane crystal by considering three-dimensional crystals formed in the potential of Eq. (55). This potential produces a crystal in a spheroidal volume of axial extent $2Z_{p}$ and diameter $2R_{p}$ with constant, three-dimensional density $n_{0}$ given by
$n_{0}=\frac{\epsilon_{0}m}{Q^{2}}\omega_{z}^{2}(1+2\beta)\,$, (56)
see Eqs. (19), (23). For this spheroidal shaped crystal, the 2D number density per unit area, $\sigma(\rho)$, is
@@
where $\sigma=n_{0}2Z_{p}$ is the 2D number density at the center. Equation (57) provides a good estimate for the 2D number density of a single-plane crystal, obtained in the limit of small $Z_{p}/R_{p}$. It is the equivalent expression for the inhomogeneity in the 2D number density for a single-plane crystal as Eq. (28) is for the inhomogeneity in the linear density for chains of ions.
@@
For finite planes, the inhomogeneous charge density of single-plane ion crystals is necessarily associated with dislocations. These are illustrated by the red dots in Fig. 22(a) [102]. Near the boundary of the planar crystal, shells are formed that conform to the overall symmetry of the confining potential [104]. Shell structure is typical for systems with long-range interactions [45]. For azimutally symmetric potentials, the shells are circular and the locations of ions in neighboring shells is staggered, see Fig. 22. This shell structure has a different symmetry than the underlying triangular lattice and contributes to the formation of dislocations.
The number of dislocations in the interior of the crystal can be reduced by modifying the confining potential through the addition of anharmonic terms, thereby approaching a regular triangular lattice. The configuration of Fig. 22(a), for instance, can be made more regular by adding a fourth-order term to the harmonic confining potential. The resulting structure is presented in Fig. 22(b) for an optimized potential and in a configuration relevant for the Penning trap. The fourth-order term is frequently called a $C_{4}$ term. Structures even closer to uniform triangular lattices can be achieved by adding corrections to the confining potential with the same underlying symmetry as the target structure. In Penning traps the symmetry of the boundary is determined by the symmetry of the rotating wall potential. Theoretical studies indicate that a nearly perfect triangular lattice can be obtained by fine tuning the strength of an $m=3$ rotating wall potential [102], where $m$ refers to the azimuthal symmetry of the applied rotating wall potential ($\propto\cos[m(\phi+\omega_{r}t)]$, see Eq. (25)). Experimental work demonstrated that this strategy leads to the formation of more regular ion crystals in Penning traps [229]. Overall, these strategies are at early stages of experimental investigation.
@@ II.2.3 Normal modes
In what follows we discuss the normal modes of single-plane ion crystals. We assume that the ions are confined to a single plane by a harmonic, conservative potential. The discussion applies to single-plane crystals in Penning traps and in linear rf traps; for the latter, micromotion is neglected and the pseudopotential approximation is assumed to be valid.
@@
We start with the normal modes of an infinite, single-plane ion crystal. This is an idealization, that provides intuition for understanding the normal modes of finite-sized, two-dimensional structures. In the infinite triangular lattice, the equilibrium positions are $\bm{\rho}=[(n_{1}+n_{2}/2)\hat{x}+(\sqrt{3}/2)\,n_{2}\,\hat{y}]a_{\mathrm{lat}}$, with $n_{1},n_{2}$ integers. The normal modes are waves with wave vector $\mathbf{k}$} in the xyx-y plane and displacements proportional to $\exp\left(i{\mathbf{k}}\cdot\bm{\rho}\right)$. For every $\mathbf{k}$} there are three different polarizations: two are in the xyx-y plane while the third is in the $\hat{z}-$direction, perpendicular to the plane. As an example consider the two in-plane modes when $k_{y}=0$. They have polarizations in the $\hat{x}$ and $\hat{y}$ directions, corresponding to compressional and transverse modes. Their dispersion relation, along with the dispersion relation for oscillations normal to the plane, is displayed in Fig. 23 for $\sigma a_{0}^{2}=1.07$ and at the instability of the single plane ($\sigma a_{0}^{2}=1.11$). Note that the bandwidth of the in-plane modes overlaps with the bandwidth of the mode for oscillations normal to the plane.
@@
The dispersion relation of the in-plane modes exhibits the characteristic features of acoustic modes in solids: the frequency increases as the wavelength decreases. The longitudinal (compressional) mode displays a sublinear dispersion at long wavelengths ($\hat{x}-$mode in Fig. 23), $\omega(k)\sim\sqrt{k}$, arising from the long-range nature of the Coulomb repulsion [338]. The dispersion relation of the out-of-plane oscillations in the $\hat{z}$ direction is similar to optical modes in solids: the mode frequency decreases as the wavelength increases. The largest frequency is the transverse confining frequency $\omega_{z}$, where the crystal oscillates as a bulk. The shortest wavelengths, comparable to the inter-ion spacing, are the smallest frequencies. As visible in Fig. 23, at the single-plane stability limit, $\left(\sigma{a_{0}}^{2}=1.1\right)$, the mode frequency becomes zero at $k_{x}a_{\mathrm{lat}}=4\pi/3$, producing a second order phase transition discussed in Sec. II.2.4.
@@
Figure 23: Normal mode frequencies of a single-plane triangular lattice as a function of $k_{x}$, in units of the inverse of the lattice constant $a_{\mathrm{lat}}=(2/\sqrt{3}\sigma)^{1/2}$. The frequencies $\omega({\mathbf{k}})$ are squared and in units of the frequency $\omega_{z}$ of the transverse harmonic trap, confining the ions on the xyx-y plane. The modes shown are the in-plane (along $\hat{x}$ and $\hat{y}$) and the transverse (along $\hat{z}$) modes for $k_{y}=0$. Dotted lines: $\sigma a_{0}^{2}=1.07$; solid lines $\sigma a_{0}^{2}=1.11$, corresponding approximately to the density $\sigma$ where the single plane becomes mechanically unstable. The branch of the transverse modes approaches 0 for $k_{x}=4\pi/3$. The phase transition of this mechanical instability is discussed in Sec. II.2.4. From 100.
@@
The normal modes of trapped, single-plane ion crystals take on the character of the infinite single-plane crystal modes, including a similar dispersion relation. The general theoretical approach consists of diagonalizing the stiffness matrix, obtained by performing the Taylor expansion of the Coulomb potential, with some salient differences between the rf- and the Penning trap due to the respective micromotion and the strong magnetic field [365, 326]. In rf traps, some of the ions in the crystal undergo significant micromotion: Including the effects of micromotion in describing normal modes is challenging and has only been incorporated in a few studies [182, 186]. In Ref. [182], including micromotion through a full time-dependent Coulomb theory was necessary to obtain good agreement between the measured and predicted eigenfrequencies for the in-plane mode spectrum of a three-ion crystal. In Ref. [186], some discrepancies with the predictions of pseudopotential theory were obtained for some of the measured transverse mode frequencies of a single-plane crystal in a rf trap composed by 8 ions. Theoretical calculations including micromotion did not resolve the discrepancies.
@@
In a Penning trap, the natural frame in which to analyze the normal modes is the frame that co-rotates with the ion crystal. Because the non-thermal ion motion is a rigid body rotation about the trap symmetry axis, it does not impact the calculation of the normal modes beyond adding a centrifugal potential. However, the strong magnetic field requires one to modify the theoretical procedure for determining the ion crystal normal modes.
@@
In general, a single-plane crystal of $N$ ions supports $2N$ in-plane modes and $N$ transverse modes. The latter are frequently called drumhead modes because they resemble the modes of a drum with open boundary conditions [310, 365, 326]. In a Penning trap, the drumhead modes for a single-plane crystal involve ion motion parallel to the magnetic field. Therefore, the magnetic field does not impact the drumhead mode frequencies or eigenvectors in the harmonic approximation, and the drumhead mode structure is the same for single-plane crystals in Penning traps and in rf traps, provided the pseudopotential approximation applies to the latter.
@@
Figure 24 illustrates the oscillation amplitudes of the four highest-frequency drumhead modes for a single-plane ion crystal. From larger to lower frequency: The highest frequency mode is the center-of-mass (or bulk) mode where all ions in the crystal oscillate as a bulk at frequency $\omega_{z}$. The next drumhead modes are tilt modes. Like the out-of-plane oscillations ($\hat{z}$ polarization in Fig. 23) in an infinite crystal, as the normal mode frequency decreases, so does the effective wavelength. In the short wave-length limit, the wavelength is of the order of the lattice spacing in the interior of the crystal and only ions in the crystal interior participate in the mode. In this spatial region the crystal approaches a regular triangular lattice and the short-wavelength drumhead modes essentially coincide with the transverse modes of the infinite crystal. See Refs. [310, 186] for examples of experimental spectroscopy of drumhead modes in single plane crystals in rf and Penning traps.
@@
Figure 24: (Color online) The four highest frequency drumhead mode eigenvectors calculated for a single-plane ion crystal consisting of $N=217$ ions. The three rows correspond to three different in-plane confinement strengths, respectively. The color scale illustrates the eigenvector amplitude with the maximal value set to one. From 365.
@@
Figure 25: Example calculation of the in-plane mode frequencies for a single-plane crystal with $N=120$ ions in the presence of no magnetic field (green central branch) and in the presence of a strong magnetic field that splits the single central branch into low-frequency (blue) and high frequency (orange) branches. The central (zero-field) branch is the in-plane mode spectrum in the pseudopotential approximation for a single-plane ion crystal in an rf trap with $\omega_{z}/(2\pi)=1.59$ MHz and effective confinements $\sqrt{\beta_{x}}\omega_{z}/(2\pi)=276$ kHz and $\sqrt{\beta_{y}}\omega_{z}/(2\pi)=259$ kHz in the orthogonal directions. The blue and orange branches are, respectively, the low-frequency ${\mathbf{E}}\times{\mathbf{B}}$ mode and high-frequency cyclotron mode spectra for the same single-plane crystal confined in a Penning trap by an identical effective potential in the rotating frame of the crystal. For the Penning trap, a cyclotron frequency $\Omega_{c}/(2\pi)=7.60$ MHz, a rotation frequency $\omega_{r}/(2\pi)=180$ kHz and a weak quadrupole rotating wall potential characterized by $\omega_{W}=68$ kHz (see Eq. (26)) are assumed. The mode frequencies are referenced to the rotating frame of the ion crystal. From 326.
@@
For single-plane crystals in an rf trap the magnetic field is nominally zero and the in-plane modes exhibit acoustic-like dispersion, similar to the in-plane modes of an infinite crystal (the $\hat{x}$ and $\hat{y}$ polarizations of Fig. 23). The green branch in Fig. 25 shows the results of a calculation of the in-plane mode spectrum for a single-plane crystal consisting of 120 ions in zero magnetic field [326]. The calculation is performed for an rf trap in the pseudopotential approximation with an axial frequency $\omega_{z}/(2\pi)=1.59$ MHz and confinement frequencies $\omega_{x}/(2\pi)=0.276$ MHz and $\omega_{y}/(2\pi)=0.259$ MHz in the orthogonal directions. For $N=120$ ions, there exist 240 in-plane modes, which are ordered in the figure according to increasing frequency. The lowest frequency modes are associated with the weak asymmetry that provides azimuthal confinement, without which there would be a zero-frequency mode corresponding to rigid body rotation about the $\hat{z}-$axis. The higher frequency modes correspond to effective shorter wavelength oscillations where ion motion becomes anti-correlated over a distance of a few inter-ion spacings.
@@
Adding a strong magnetic field normal to the single-plane crystal (parallel to the $z$-axis) gives rise to velocity dependent forces. This is the case of the Penning trap. Here, the mode analysis becomes more involved for the in-plane modes, requiring the solution of a generalized eigenvalue problem [365, 103, 326]. Figure 25 illustrates that an important consequence of the strong magnetic field is to split the single band of $2N$ in-plane modes for $B=0$ (green central branch) into a band of $N$ high frequency cyclotron modes and a band of $N$ low frequency ${\mathbf{E}}\times{\mathbf{B}}$ modes with frequency on the order of the single ion magnetron frequency $\omega_{m}$ (see Sec. I.2.2).